Solve: 2x(squared)-7x=-5
Answer by unlockmath(1688)

Hello,
2x^2-7x=-5
First, add 5 to both sides.
2x^2-7x+5=0
This can be factored as:
(2x-5)(x-1)=0
Solve for x to get:
x=5/2
x=1
There we go. Got it?
RJ
www.math-unlock.com